The quantity p-q is a factor of how many of the following?
p(squared)- q(squared) p(squared)+ q(squared) p(cubed)-q(cubed)
p(cubed)+q(cubed)
A.none B. only one C.two only D.three only E.Four
: The quantity p-q is a factor of how many of the following?
p(squared)- q(squared) p(squared)+ q(squared) p(cubed)-q(cubed)
p(cubed)+q(cubed)
A.none B. only one C.two only D.three only E.Four

Answer by Earlsdon(3816)

The quantity (p-q) is a factor of?
1) p^2-q^2 = highlight((p-q))(p+q)
2) p^2+q^2 Prime!
3) p^3-q^3 = highlight((p-q))(p^2+pq+p^2)
4) p^3+q^3 = (p+q)(p^2-pq+q^2)
Answer C is appropriate!